Find the general solution to 13 - 02 X o 5 203
Answer #1

First find eign value of matrix then eign vector corresponding to eign value then put into general form solution.Page 1 8 Pind general Solukon to 3 0 -2 0 3 2 02 Solution 3 -2 frost find eign value of matrix A= S wo 2 Andito Lohore A is cfor x=3+24 page-2 - 2 3-(3+21) o 6-3+2) 92 0 13 poo 2 3-(3+21) -2i -2 al 92 > 2-2 D -2 03 0 293 oh=-29372 -zéu = 293=0 (2-2)= (ي=يا ا | Page 3 وله كما ۷( = a2) = D VGA-3+4)= & generel Solukoo y = quietit + care erzt + cauzelat. = لا C st e + C (3-2)